TUGboat, Volume 0 (2060), No. 0 1001

Example TUGboat Regular Article

First Last

Abstract

This is an example article for a regular TUGboat issue.

1 Introduction

This is an example article for TUGboat, from http:

//tug.org/TUGboat/location.html.

⋄ First Last Street Address Town, Postal Country

                 user@example.org
                 http://example.org/~user/